A ball was thrown in the air and I have the formula
h(t)=-4.9t(squared)+31t+1
31 is the initial velocity and 1 (meter) is the initail height
i need to find the height after 1 second. i also need to find the time it is at ten feet.
Do i use the quadradic formula and put the numbers in? if so, where do i put them in?

Answer by ankor@dixie-net.com(22740)

A ball was thrown in the air and I have the formula
h(t)=-4.9t^2 + 31t + 1
:
find the height after 1 second.
Replace t with 1 in the above equation
h(t) = -4.9(1^2) + 31(1) + 1
h(t) = -4.9 + 31 + 1
h(t) = 27.1 m after 1 sec
:
find the time it is at ten feet, I'm going to assume you mean 10 meters here
-4.9t^2 + 31t + 1 = h(t)
replace h(t) with 10, find t
-4.9t^2 + 31t + 1 = 10
:
-4.9t^2 + 31t + 1 - 10 = 0
:
-4.9t^2 + 31t - 9 = 0
Use the quadratic formula to find t
a=-4.9; b=31; c=-9
:
You should get two solutions, t=.3 sec and t=6.05 sec
one time when it's at 10m on the way up and, another solution for 10m on the way down